Aluminum bromide is produced by the reaction of aluminum metal with bromine liquid, according to the following balanced equation:

2 Al(s) + 3 Br2(l) → 2 AlBr3(aq)

If 10.0 grams of Al is added to 10.0 g of Br2, assuming the reaction has yielded 100% yield, calculate the mass (in grams) of AlBr3 produced in this reaction
1
Expert's answer
2020-05-15T10:23:14-0400

Solution.

"2Al + 3Br2 = 2AlBr3"

"n(Al) = \frac{10}{26.98} = 0.37 \ mol"

"n(Br2) = \frac{10}{159.80}=0.06 \ mol"

In this reaction, Br2 is limiting reagent

"n(AlBr3) = \frac{0.06 \times 2}{3} = 0.04 \ mol"

Because yield of the reaction is 100 %, that:

"m(AlBr3) = 0.04 \times 266.68 = 10.67 \ g"

Answer:

m(AlBr3) = 10.67 g